which of the following is a linear equation in x;y and z?


1. x-e^2y=3z, where e=2.71828....


2. 2πIn(e 1/z)-2y+z=In(3)-x


3. √y^-2 +4y-2z=7x


4. y+4y-2z=7x^3


1
Expert's answer
2021-05-17T02:32:02-0400

The equation "x-e^2 y=3z" is a linear equation in x, y and z because-

      

   Since, "x-e^2y-3z=0"


here, "a=1,b=-e^2,c=-3"

      Here "a\neq b\neq c\neq 0"

 

   SO It is a linear equation in x, y and z.